# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v1

@@ 4. Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)
Lemma 4.1 (Transformation of the blanket divergence for FFT).
@@
The proof is provided in Appendix A.5. Compared with the transformations in Lemma 3.1, here we work with a representation of the blanket divergence that does not involve explicit expectations or integrals, which is particularly convenient for deriving rigorous bounds on the truncation error. In particular, in our FFT-based computation we must truncate the support of the random variable in order to approximate an a priori unbounded distribution on a finite grid. This issue is especially relevant for non-pure local DP (see, e.g., the discussion of the Gaussian mechanism in Balle et al. (Balle et al., 2019)); hence this formulation is crucial for our purposes.
Su et al. (Su et al., 2025) also propose an FFT-based accountant for the blanket divergence, but their running time scales as $O(n^{2})$ in the number of users when targeting a constant additive error. In contrast, by adopting the concentration-inequality-based error control of Gopi et al. (Gopi et al., 2021) with our results in Section 3, we can tune the truncation and discretization parameters so that our algorithm runs in near-linear time $\widetilde{O}(n)$ while achieving a prescribed relative error. Algorithms 1 and 2 together form our FFT-based procedure.
4.1. FFT Algorithm
@@ 5. Experiments
Figure 1. Experimental validation of the FFT-based blanket-divergence accountant and the shuffle-index analysis.
Four labeled subfigures: two plots illustrating the accuracy–runtime trade-offs of the FFT-based accountant, one combined plot showing the tightness of the shuffle-index-based bands for 3-$R$ and Gaussian mechanisms, and one plot showing the dependence of the lower shuffle index on the effective per-user standard deviation.
@@
5.1. Privacy Curves and Shuffle Index
@@
5.2. Accuracy and Runtime of the FFT Accountant
@@ 6. Limitations and Conclusion
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
@@ Appendix A Missing Proofs
(4) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{n})_{+}\bigr],\qquad(S_{n})_{+}:=\max\{S_{n},0\}.$
@@
(5) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{\gamma}\,\mathbb{E}\bigl[Z_{1}\,\mathbf{1}\{S_{n}>0\}\bigr].$
@@
(6) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{\gamma}\,\mathbb{E}\bigl[Z_{1}\,\Pr\bigl[S_{n-1}>-Z_{1}\bigr]\bigr].$
@@
(7) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{n\gamma}\int_{0}^{\infty}\Pr\left[S_{n}>u\right]\,du.$
@@
Theorem A.1 (Triangular-array Edgeworth expansion after Angst–Poly (Angst and Poly, 2017)).
@@
Theorem A.3 (Triangular-array moderate deviations after Slastnikov (Slastnikov, 1979)).
@@
Lemma A.6 (Truncation error for the blanket divergence).
@@
(8) $\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\;\leq\;\frac{C}{\tau_{n}^{p-1}},\qquad n\geq 1.$
@@
(9) $\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\leq\frac{1}{n\gamma}\,\mathbb{E}\biggl|\sum_{i=1}^{n}\Delta_{i,n}\biggr|.$
@@
Lemma A.7 (Evaluation of the expectation).
@@
(10) $\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}\bigl(t_{n}-\eta_{n}(Z)\bigr)\bigr)\Bigr]=\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+r_{n}\bigr),$
@@
(11) $\sup_{u\in\mathbb{R}}\biggl|F_{n-1}(u)-\Phi(u)-\sum_{j=1}^{s-2}(n-1)^{-j/2}P_{j,n-1}(u)\varphi(u)\biggr|\leq C_{s}(n-1)^{-(s-2)/2}.$
@@
(12) $1-\Phi\bigl(t_{n}-\eta_{n}(z)\bigr)=g(t_{n})-g^{\prime}(t_{n})\eta_{n}(z)+\frac{1}{2}g^{\prime\prime}(\xi)\eta_{n}(z)^{2}.$
@@
(13) $A_{n}=\mu_{n}(1-\Phi(t_{n}))+\frac{\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\varphi(t_{n})+O\Bigl(\frac{\varphi(t_{n})(1+t_{n})}{n-1}\Bigr).$
@@
(14) $\quad+O\Bigl(\frac{\varphi(t_{n})t_{n}^{4}}{\sigma_{\varepsilon_{n}}^{2}(n-1)}\mathbb{E}[Z^{3}]\Bigr).$
@@
(15) $B_{n,1}=\frac{\kappa_{3}\mu_{n}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\varphi(t_{n})(t_{n}^{2}-1)+\frac{\kappa_{3}(\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2})}{6\sigma_{\varepsilon_{n}}^{4}}\varphi(t_{n})t_{n}(t_{n}^{2}-3)\frac{1}{n-1}+O\Bigl(\frac{\varphi(t_{n})(1+t_{n}^{4})}{(n-1)^{3/2}}\Bigr).$
@@
(16) $\leq C_{j}^{\prime\prime}(n-1)^{-j/2}\,\varphi(t_{n})\,(1+t_{n}^{k_{j}}),$
@@
(17) $\quad+O\bigl(\tau_{n}(n-1)^{-(s-2)/2}\bigr).$
@@
(18) $\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}(t_{n}-\eta_{n}(Z))\bigr)\Bigr]=\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+r_{n}\bigr),$
@@
(19) $\sigma_{\varepsilon_{n}}=\sqrt{\gamma}\,\sigma+O(\varepsilon_{n}).$
@@
(20) $\sigma_{\varepsilon_{n}}^{3}=\sigma_{0}^{3}\bigl(1+o(1)\bigr)=\bigl(\sqrt{\gamma}\,\sigma\bigr)^{3}\bigl(1+o(1)\bigr)=\gamma^{3/2}\sigma^{3}\bigl(1+o(1)\bigr).$
@@
Corollary A.8 (Discrete bounded case).
@@
(21) $\mathcal{D}^{\mathrm{blanket}}_{e^{{\varepsilon_{n}}},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{\sigma_{\varepsilon_{n}}}{\sqrt{n}\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$
@@
(22) $\sup_{0\leq x\leq c\sqrt{\log n}}\left|\frac{\Pr\left[T_{n}>x\right]}{1-\Phi(x)}-1\right|\leq\Delta_{n},\qquad n\to\infty.$
@@
(23) $\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\,\bigl(\mathrm{Main}_{n}+\mathrm{Err}_{n}\bigr).$
@@
(24) $\mathrm{Main}_{n}=\int_{t_{n}}^{\infty}(1-\Phi(x))\,dx=\frac{\varphi(t_{n})}{t_{n}^{2}}\bigl(1+O(t_{n}^{-2})\bigr),\qquad t_{n}\to\infty.$
@@
(25) $\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\,\frac{\varphi(t_{n})}{t_{n}^{2}}\,\bigl(1+o(1)\bigr),\qquad t_{n}:=-\frac{\mu_{n}\sqrt{n}}{\sigma_{\varepsilon_{n}}}.$
@@
(26) $\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;g_{n}\bigl(\varepsilon_{n};\chi(x_{1},x_{1}^{\prime};\mathcal{R}_{\mathrm{ref}})\bigr)\,\bigl(1+\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr),\qquad\bigl|\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr|\leq\rho_{n},$
@@
Lemma A.9 (Pairwise ordering and equality condition).
@@
(27) $\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}=\mathbb{E}_{\begin{subarray}{c}M\sim 1+\mathrm{Bin}(n-1,\gamma)\\ Y_{1:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\Bigl[l_{\varepsilon}(Y_{1})\,\mathbf{1}\Bigl\{\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr\}\Bigr].$
@@ A.6. The proof of Theorem 5
This proof is an adaptation of the FFT error analysis of Gopi et al. (Gopi et al., 2021) to our blanket-divergence setting.
@@
(28) $p(x):=\Pr_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[l_{\varepsilon}(Y_{1};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})+\sum_{i=2}^{M}l_{\varepsilon}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})>0\Bigr].$
@@
(29) $\leq\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]+\Pr\bigl[\perp\bigr],$
(30) $p(x)$ $\geq\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]-\Pr\bigl[\perp\bigr].$
@@
(31) $\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]\;\leq\;E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>-c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$
@@
(32) $\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]\;\geq\;-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>+c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$
@@
(33) $=E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+\sum_{z}\Pr[S=z]\,\Pr\bigl[l(Y_{1})>-c-z-\mathbb{E}[\textstyle\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}]\bigr],$
@@
(34) $p(x)$ $\leq\widehat{p}_{c}(x)+E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+E_{\mathrm{trunc}}(x).$
@@
(35) $p(x)$ $\geq\widehat{p}_{-c}(x)-E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})-E_{\mathrm{trunc}}(x).$
@@
(36) $p(x)$ $\leq P(c,F_{x})+E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+2E_{\mathrm{trunc}}(x),$
@@
(37) $p(x)$ $\geq P(-c,F_{x})-E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})-2E_{\mathrm{trunc}}(x).$
@@
(38) $\leq\frac{1}{2}\bigl|P(c;x_{1})-P(-c;x_{1})\bigr|+\frac{\mathrm{e}^{\varepsilon}}{2}\bigl|P(c;x_{1}^{\prime})-P(-c;x_{1}^{\prime})\bigr|+\frac{\delta^{\mathrm{up}}_{\mathrm{err}}+\delta^{\mathrm{low}}_{\mathrm{err}}}{2}.$
@@
(39) $P(c;x)-P(-c;x)$ $=(1-\Pr[\perp])\,\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[-c<\widetilde{S}_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}\leq c\,\big|\,Z_{1}\bigr]\Bigr],$
@@
(40) $\qquad\leq\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]+\Pr\bigl[|E_{n-1}|>c/2\bigr].$
@@
(41) $|P(c;x)-P(-c;x)|$ $\leq\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]\Bigr]+\Pr\bigl[|E_{n-1}|>c/2\bigr].$
@@
(42) $\Pr\bigl[|E_{n-1}|>c/2\bigr]\leq\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c/6\bigr]+\Pr\bigl[|E_{\mathrm{Mdisc}}|>c/6\bigr]+\Pr\bigl[|E_{\mathrm{Malias}}|>c/6\bigr].$
@@
(43) $\quad+\frac{1+\mathrm{e}^{\varepsilon}}{2}\,\Pr\bigl[|E_{\mathrm{Malias}}|>c/6\bigr]+\delta_{\mathrm{alias}}.$

### v2

@@ 4 Analysis of Blanket Divergence via the Fast Fourier Transform (FFT)
Remark 4.1 (Transformation of the blanket divergence for FFT).
@@
The proof is provided in Appendix A.5. Compared with the transformations in Lemma 3.1, here we work with a representation of the blanket divergence that does not involve explicit expectations or integrals, which is particularly convenient for deriving rigorous bounds on the truncation error. In particular, in our FFT-based computation we must truncate the support of the random variable in order to approximate an a priori unbounded distribution on a finite grid. This issue is especially relevant for non-pure local DP (see, e.g., the discussion of the Gaussian mechanism in Balle et al. Balle et al. (2019)); hence this formulation is crucial for our purposes.
Su et al. Su et al. (2025) also propose an FFT-based accountant for the blanket divergence, but their running time scales as $O(n^{2})$ in the number of users when targeting a constant additive error. In contrast, by adopting the concentration-inequality-based error control of Gopi et al. Gopi et al. (2021) with our results in Section 3, we can tune the truncation and discretization parameters so that our algorithm runs in near-linear time $\widetilde{O}(n)$ while achieving a prescribed relative error. Algorithms 1 and 2 together form our FFT-based procedure.
4.1 FFT Algorithm
@@ 5 Experiments
Figure 1: Experimental validation of the FFT-based blanket-divergence accountant and the shuffle-index analysis.
@@
5.1 Privacy Curves and Shuffle Index
@@
5.2 Accuracy and Runtime of the FFT Accountant
@@ Appendix A Missing Proofs
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{n\gamma}\,\mathbb{E}\bigl[(S_{n})_{+}\bigr],\qquad(S_{n})_{+}:=\max\{S_{n},0\}.$ (4)
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{\gamma}\,\mathbb{E}\bigl[Z_{1}\,\mathbf{1}\{S_{n}>0\}\bigr].$ (5)
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{\gamma}\,\mathbb{E}\bigl[Z_{1}\,\Pr\bigl[S_{n-1}>-Z_{1}\bigr]\bigr].$ (6)
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{1}{n\gamma}\int_{0}^{\infty}\Pr\left[S_{n}>u\right]\,du.$ (7)
@@
Theorem A.1 (Triangular-array Edgeworth expansion after Angst–Poly Angst and Poly (2017)).
@@
Theorem A.3 (Triangular-array moderate deviations after Slastnikov Slastnikov (1979)).
@@
Remark A.5 (Truncation error for the blanket divergence).
@@
$\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\;\leq\;\frac{C}{\tau_{n}^{p-1}},\qquad n\geq 1.$ (8)
@@
$\bigl|\mathcal{D}^{\mathrm{blanket}}-\mathcal{D}^{\mathrm{tblanket}}\bigr|\leq\frac{1}{n\gamma}\,\mathbb{E}\biggl|\sum_{i=1}^{n}\Delta_{i,n}\biggr|.$ (9)
@@
Remark A.6 (Evaluation of the expectation).
@@
$\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}\bigl(t_{n}-\eta_{n}(Z)\bigr)\bigr)\Bigr]=\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+r_{n}\bigr),$ (10)
@@
$\sup_{u\in\mathbb{R}}\biggl|F_{n-1}(u)-\Phi(u)-\sum_{j=1}^{s-2}(n-1)^{-j/2}P_{j,n-1}(u)\varphi(u)\biggr|\leq C_{s}(n-1)^{-(s-2)/2}.$ (11)
@@
$1-\Phi\bigl(t_{n}-\eta_{n}(z)\bigr)=g(t_{n})-g^{\prime}(t_{n})\eta_{n}(z)+\frac{1}{2}g^{\prime\prime}(\xi)\eta_{n}(z)^{2}.$ (12)
@@
$A_{n}=\mu_{n}(1-\Phi(t_{n}))+\frac{\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\varphi(t_{n})+O\Bigl(\frac{\varphi(t_{n})(1+t_{n})}{n-1}\Bigr).$ (13)
@@
$\quad+O\Bigl(\frac{\varphi(t_{n})t_{n}^{4}}{\sigma_{\varepsilon_{n}}^{2}(n-1)}\mathbb{E}[Z^{3}]\Bigr).$ (14)
@@
$B_{n,1}=\frac{\kappa_{3}\mu_{n}}{6\sigma_{\varepsilon_{n}}^{3}\sqrt{n-1}}\varphi(t_{n})(t_{n}^{2}-1)+\frac{\kappa_{3}(\sigma_{\varepsilon_{n}}^{2}+\mu_{n}^{2})}{6\sigma_{\varepsilon_{n}}^{4}}\varphi(t_{n})t_{n}(t_{n}^{2}-3)\frac{1}{n-1}+O\Bigl(\frac{\varphi(t_{n})(1+t_{n}^{4})}{(n-1)^{3/2}}\Bigr).$ (15)
@@
$\leq C_{j}^{\prime\prime}(n-1)^{-j/2}\,\varphi(t_{n})\,(1+t_{n}^{k_{j}}),$ (16)
@@
$\quad+O\bigl(\tau_{n}(n-1)^{-(s-2)/2}\bigr).$ (17)
@@
$\mathbb{E}\Bigl[Z\bigl(1-F_{n-1}(t_{n}-\eta_{n}(Z))\bigr)\Bigr]=\frac{\varphi(t_{n})}{\sqrt{n-1}}\,\frac{\sigma_{\varepsilon_{n}}}{t_{n}^{2}}\,\bigl(1+r_{n}\bigr),$ (18)
@@
$\sigma_{\varepsilon_{n}}=\sqrt{\gamma}\,\sigma+O(\varepsilon_{n}).$ (19)
@@
$\sigma_{\varepsilon_{n}}^{3}=\sigma_{0}^{3}\bigl(1+o(1)\bigr)=\bigl(\sqrt{\gamma}\,\sigma\bigr)^{3}\bigl(1+o(1)\bigr)=\gamma^{3/2}\sigma^{3}\bigl(1+o(1)\bigr).$ (20)
@@
Remark A.7 (Discrete bounded case).
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{{\varepsilon_{n}}},n,\mathcal{R}_{\mathrm{ref}},\gamma}(\mathcal{R}_{x_{1}}\|\mathcal{R}_{x_{1}^{\prime}})=\frac{\sigma_{\varepsilon_{n}}}{\sqrt{n}\gamma}\,\int_{t_{n}}^{\infty}\bigl(1-F_{n}(x)\bigr)\,dx.$ (21)
@@
$\sup_{0\leq x\leq c\sqrt{\log n}}\left|\frac{\Pr\left[T_{n}>x\right]}{1-\Phi(x)}-1\right|\leq\Delta_{n},\qquad n\to\infty.$ (22)
@@
$\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\,\bigl(\mathrm{Main}_{n}+\mathrm{Err}_{n}\bigr).$ (23)
@@
$\mathrm{Main}_{n}=\int_{t_{n}}^{\infty}(1-\Phi(x))\,dx=\frac{\varphi(t_{n})}{t_{n}^{2}}\bigl(1+O(t_{n}^{-2})\bigr),\qquad t_{n}\to\infty.$ (24)
@@
$\mathcal{D}^{\mathrm{blanket}}=\frac{\sigma_{\varepsilon_{n}}}{\gamma\sqrt{n}}\,\frac{\varphi(t_{n})}{t_{n}^{2}}\,\bigl(1+o(1)\bigr),\qquad t_{n}:=-\frac{\mu_{n}\sqrt{n}}{\sigma_{\varepsilon_{n}}}.$ (25)
@@
$\mathcal{D}^{\mathrm{blanket}}_{e^{\varepsilon_{n}},n,\mathcal{R}_{\mathrm{ref}},\gamma}\;=\;g_{n}\bigl(\varepsilon_{n};\chi(x_{1},x_{1}^{\prime};\mathcal{R}_{\mathrm{ref}})\bigr)\,\bigl(1+\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr),\qquad\bigl|\theta_{n}(x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})\bigr|\leq\rho_{n},$ (26)
@@
Remark A.8 (Pairwise ordering and equality condition).
@@
$\mathcal{D}_{\mathrm{e}^{\varepsilon},n,\mathcal{R}_{\mathrm{ref}},\gamma}^{\text{blanket}}=\mathbb{E}_{\begin{subarray}{c}M\sim 1+\mathrm{Bin}(n-1,\gamma)\\ Y_{1:n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{R}_{\mathrm{ref}}\end{subarray}}\Bigl[l_{\varepsilon}(Y_{1})\,\mathbf{1}\Bigl\{\sum_{i=1}^{M}l_{\varepsilon}(Y_{i})>0\Bigr\}\Bigr].$ (27)
@@ A.6 The proof of Theorem 5
This proof is an adaptation of the FFT error analysis of Gopi et al. Gopi et al. (2021) to our blanket-divergence setting.
@@
$p(x):=\Pr_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[l_{\varepsilon}(Y_{1};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})+\sum_{i=2}^{M}l_{\varepsilon}(Y_{i};x_{1},x_{1}^{\prime},\mathcal{R}_{\mathrm{ref}})>0\Bigr].$ (28)
@@
$\leq\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]+\Pr\bigl[\perp\bigr],$ (29)
$p(x)$ $\geq\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]-\Pr\bigl[\perp\bigr].$ (30)
@@
$\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]\;\leq\;E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>-c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$ (31)
@@
$\Pr\Bigl[l(Y_{1})+\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}>0\Bigr]\;\geq\;-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+\Pr\Bigl[\sum_{i=2}^{M}Z_{i}^{\mathrm{di}}>+c-l(Y_{1})-\Delta_{\mathrm{ES}}\Bigr].$ (32)
@@
$=E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+\sum_{z}\Pr[S=z]\,\Pr\bigl[l(Y_{1})>-c-z-\mathbb{E}[\textstyle\sum_{i=2}^{M}Z_{i}^{\mathrm{tr}}]\bigr],$ (33)
@@
$p(x)$ $\leq\widehat{p}_{c}(x)+E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+E_{\mathrm{trunc}}(x).$ (34)
@@
$p(x)$ $\geq\widehat{p}_{-c}(x)-E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})-E_{\mathrm{trunc}}(x).$ (35)
@@
$p(x)$ $\leq P(c,F_{x})+E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})+E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})+2E_{\mathrm{trunc}}(x),$ (36)
@@
$p(x)$ $\geq P(-c,F_{x})-E_{\mathrm{alias}}(x;h,s,w^{\mathrm{in}},w^{\mathrm{out}})-E_{\mathrm{disc}}(x;c,h,s,w^{\mathrm{in}})-2E_{\mathrm{trunc}}(x).$ (37)
@@
$\leq\frac{1}{2}\bigl|P(c;x_{1})-P(-c;x_{1})\bigr|+\frac{\mathrm{e}^{\varepsilon}}{2}\bigl|P(c;x_{1}^{\prime})-P(-c;x_{1}^{\prime})\bigr|+\frac{\delta^{\mathrm{up}}_{\mathrm{err}}+\delta^{\mathrm{low}}_{\mathrm{err}}}{2}.$ (38)
@@
$P(c;x)-P(-c;x)$ $=(1-\Pr[\perp])\,\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[-c<\widetilde{S}_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}\leq c\,\big|\,Z_{1}\bigr]\Bigr],$ (39)
@@
$\qquad\leq\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]+\Pr\bigl[|E_{n-1}|>c/2\bigr].$ (40)
@@
$|P(c;x)-P(-c;x)|$ $\leq\mathbb{E}_{Y_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c\,\big|\,Z_{1}\bigr]\Bigr]+\Pr\bigl[|E_{n-1}|>c/2\bigr].$ (41)
@@
$\Pr\bigl[|E_{n-1}|>c/2\bigr]\leq\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c/6\bigr]+\Pr\bigl[|E_{\mathrm{Mdisc}}|>c/6\bigr]+\Pr\bigl[|E_{\mathrm{Malias}}|>c/6\bigr].$ (42)
@@
$\quad+\frac{1+\mathrm{e}^{\varepsilon}}{2}\,\Pr\bigl[|E_{\mathrm{Malias}}|>c/6\bigr]+\delta_{\mathrm{alias}}.$ (43)